OpenStudy (anonymous):

Determine algebraically whether the function is even, odd, or neither even nor odd. f(x) = -9 x3 + 8x

OpenStudy (mathmale):

Although I can understand what you're trying to convey here, it'd be a lot better if you use the correct symbols to express the power of a quantity. Instead of f(x) = -9 x3 + 8x , please write f(x)=-9*x^3 + 8x.

OpenStudy (anonymous):

-9x^3 + 8x

OpenStudy (mathmale):

Look at the two powers of x. One is x^3, the other is just x^1, or simply x. f(x)=f(x)=-9x^3+8x^1

OpenStudy (mathmale):

Are those powers of x even or odd?

OpenStudy (anonymous):

@mathmale odd?